# Generative Refocusing: Flexible Defocus Control from a Single Image

### v2

@@ Abstract
Depth-of-field control is essential in photography, but getting the perfect focus often takes several tries or special equipment. Single-image refocusing is still difficult. It involves recovering sharp content and creating realistic bokeh. Current methods have significant drawbacks. They need all-in-focus inputs, depend on synthetic data from simulators, and have limited control over aperture. We introduce Generative Refocusing, a two-step process that uses DeblurNet to recover all-in-focus images from various inputs and BokehNet for creating controllable bokeh. Our main innovation is semi-supervised training. This method combines synthetic paired data with unpaired real bokeh images, using EXIF metadata to capture real optical characteristics beyond what simulators can provide. Our experiments show we achieve top performance in defocus deblurring, bokeh synthesis, and refocusing benchmarks. Additionally, our Generative Refocusing allows text-guided adjustments and custom aperture shapes. Project page: https://generative-refocusing.github.io/
@@ 1 Introduction
Controlling depth of field and focus is essential for artistic photography. It helps guide the viewer’s eye through selective focus and bokeh effects. Getting perfect focus often requires multiple shots or special equipment, which can be tough for everyday photographers. Single-image refocusing removes these obstacles. It allows for focus and bokeh adjustments after the photo is taken. The main challenge is recovering sharp details from blurred areas while creating natural-looking bokeh, while still giving users fine control.
@@
Existing research addresses different aspects of this problem. Defocus deblurring methods [1, 8, 36, 56, 67, 73, 84] aim to recover sharp images. Diffusion approaches [41, 42] show promise for addressing spatially variant blur. Bokeh synthesis methods [9, 30, 48, 60] focus on rendering realistic depth-of-field. Recent efforts include physics-based [38, 37, 50], neural [48, 49], and diffusion-based [19, 82, 71] approaches. However, most methods handle either deblurring or bokeh synthesis individually. End-to-end refocusing often requires specialized capture setups [40, 46, 2].
@@
Despite these advances, three main limitations remain (Tab. 1). First, most methods rely on all-in-focus inputs and accurate depth maps. This restricts their use to images that already have defocus blur. Second, while synthetic data [48, 89] can be used for training, it is limited by the realism of the simulators. This approach often fails to capture detailed visual appearances. On the other hand, obtaining real paired datasets [60] is very expensive. Third, current methods usually only support aperture size, not its shape. This restriction limits the possibilities for creative bokeh effects.
@@
Table 1: Comparison across methods.  = supported;  = not $\LARGE\triangle = supported$ with caveat. In the synthetic/real training data columns, “paired” indicates the use of paired supervision, while “unpaired” indicates real unpaired data (no real paired supervision). Bokehlicious typically assumes an all-in-focus input and requires fine-tuning for different input types. For BokehMe, only the classical renderer variant supports aperture-shape control. UW = ultra-wide, W = wide.
@@
Method Input Training data Supported user control
@@
Setup AIF required? Synthetic Real Aperture size Aperture shape Focus point Defocus map
[…]
DC2 [2] Dual-view (UW+W) paired
Bokehlicious [60] Single-view paired
BokehDiff [89] Single-view paired
DiffCamera [71] Single-view paired
Ours Single-view paired unpaired
@@
To tackle these issues, we present Generative Refocusing, a flexible single-image refocusing system built on a two-stage design. Our DeblurNet module produces a sharp, in-focus image from a variety of inputs. It uses a diffusion model guided by initial deblurring predictions. Our BokehNet creates fully customizable bokeh, considering user-defined focus planes, bokeh intensity, and aperture shapes (LABEL:fig:teaser). A significant advancement in our method is our training scheme, which makes use of both 1) paired synthetic data to maintain geometric consistency and 2) unpaired real-world bokeh images with camera metadata that help us learn real lens characteristics not captured by simulators. Our approach achieves top performance in defocus deblurring, bokeh generation, and refocusing, while maintaining the natural consistency of scenes.
@@
  • A flexible refocusing pipeline accepts images in any focus state. It gives users control over the focus plane, bokeh intensity, and aperture shape through two-stage decomposition.
  • This semi-supervised framework connects synthetic and unpaired real bokeh images. It enables genuine optical learning that goes beyond what a simulator can offer. This is achieved through parameter estimation from EXIF metadata.
  • The system performs well across defocus deblurring, bokeh synthesis, and refocusing. This has been validated on existing benchmarks (DPDD [1], RealDOF [36]) and our new light-field datasets (LF-Bokeh, LF-Refocus). It also has applications in text-guided deblurring and creating aperture shapes.
@@ 2 Related Work
Figure 2: Pipeline Overview. Our method decomposes single-image refocusing into two stages: (a) Defocus Deblurring and (b) Shallow Depth-of-Field Synthesis. Given a blurry input image $I_{\text{in}}$, we optionally apply a pre-deblurring method [56] to obtain a conservative estimate $I_{\text{syn}}$, then feed both $I_{\text{in}}$ and $I_{\text{syn}}$ into DeblurNet to recover a high-quality all-in-focus image $I_{\text{aif}}$. The VAE encoder $\mathcal{E}$ and decoder $\mathcal{D}$ convert images to latent representations for processing by the DiT backbone. In the second stage, BokehNet takes the all-in-focus image $I_{\text{aif}}$, the defocus map $D_{\text{def}}$, and optionally a specific aperture shape as inputs to synthesize the refocused output. The defocus map $D_{\text{def}}$ is computed based on the estimated depth map $D$ [7], along with both the user-specified focus plane $S_{1}$ and bokeh level $K$.
Diffusion Models for Image Restoration.
Diffusion models [24, 65, 81, 68] advance image restoration via generative priors, extending to zero-shot restoration and artifact removal. Methods evolved from pixel-space [57] to efficient latent-space approaches [54, 76], enabling faster inference. ResShift [83] cuts steps from 1000 to 4–15 via residual shifting. Deblurring progressed from merging diffusion priors with regression tasks, such as matting [74] or hierarchical restoration [13], to learning spatially-varying kernels in latent space [33]. Cascaded pipelines [25] refine at multiple resolutions, while principled frameworks [45, 44] treat degradations as stochastic processes. Training-free methods [91, 32] enable plug-and-play restoration. Unlike these general methods, we target spatially-varying defocus blur with a two-stage pipeline explicitly separating deblurring from bokeh synthesis for controllable refocusing.
Defocus Deblurring.
@@
Bokeh Rendering.
Bokeh rendering progressed from physically based scattering [50, 38, 37] and differentiable rendering [63] to neural methods combining ray tracing with learned enhancements [48, 49]. Learning-based approaches evolved from fixed apertures [30] to variable f-stops [60, 9], extending to video [80] and 3D scenes [62, 72]. Diffusion models enabled bokeh control [19, 82], mainly for text-to-image synthesis. These methods require all-in-focus inputs and accurate depth. We remove this constraint via flexible input handling, which accepts defocused or all-in-focus images, and semi-supervised learning on unpaired real-world bokeh to capture authentic lens characteristics beyond simulators.
Single-Image Refocusing.
Post-capture refocusing manipulates focus after capture. Light-field cameras [46, 40] enable optical refocusing but need specialized capture. Computational approaches evolved from deconvolution [4, 86] producing artifacts, to GAN-based paradigms [58], to diffusion methods [71]. 3D representations [62, 72, 35, 43] achieve real-time controllable DoF but need multi-view capture or calibration. We combine three capabilities: (1) semi-supervised learning mixing synthetic and unpaired real data for authentic optics, (2) single-image flexibility accepting any input without preprocessing, and (3) comprehensive control over focus, intensity, and aperture shape.
Camera-Conditioned Diffusion.
Conditioning diffusion models on camera parameters progressed from extrinsic control for view synthesis [26, 14] and video generation [22, 75], to intrinsic control of focal length, aperture, and exposure [18, 12]. Direct conditioning often yields inconsistent content. Recent work uses temporal modeling [82], explicit calibration [6, 16], or transformer-based pose handling [3, 59]. These focus on generation; we operate in editing, manipulating existing images with scene consistency. By decoupling into deblurring and bokeh stages, we achieve efficient single-image manipulation without multi-frame overhead.
Semi-Supervised Learning.
When paired data is scarce, semi-supervised learning provides alternatives. The field evolved from learning without clean targets [39, 34, 27] to pseudo-labeling [11, 23, 69]. Foundation models enabled the generation of high-quality pseudo-labels [21, 85]. Unpaired frameworks [31, 90, 64] learn enhancement without paired supervision. Two-stage approaches [29, 78] bridge fidelity and perceptual quality via different supervision types. We adapt these to bokeh rendering: perfect paired refocusing data is nearly impossible to obtain. We train on synthetic pairs for structure, then unpaired real bokeh for realism. Deblur restores to all-in-focus, enabling bokeh to learn authentic lens characteristics exceeding simulators.
@@ 3 Method
Single-image refocusing couples two operations: recovering sharp content in out-of-focus regions and applying controllable bokeh to originally sharp areas. Blur magnitude is dictated by a defocus map parameterized by a user-specified focus point and the scene depth. When the input is blurry, monocular depth estimation is brittle, undermining precise control over the defocus map and often yielding misfocus or artifacts. We propose Generative Refocusing, which decomposes the task into defocus deblurring and bokeh synthesis, and comprises two models, DeblurNet and BokehNet, that jointly enable precise and controllable refocusing (see Fig. 2 for an overview of the pipeline); Such a flexible pipeline allows us to exert greater control over input conditions (see Tab. 1).
Figure 3: Training data generation. Each training sample consists of the following five components: (i) a bokeh image, (ii) an all-in-focus (AIF) image, (iii) a depth map $D$, (iv) a bokeh level $K$, and (v) a focus plane $S_{1}$. We construct these samples via three routes: (a) Synthetic paired data. Given real AIF images and depth maps $D$, we compute a defocus map $D_{\text{def}}$ parameterized by a specified bokeh level $K$ and focus plane $S_{1}$, and feed it into a bokeh renderer [48] to synthesize corresponding bokeh images. (b) Real unpaired data. Given real bokeh images, DeblurNet recovers an AIF image. We then estimate depth and extract a foreground mask [88] to define the focus plane $S_{1}$. The bokeh level $K$ is computed from the EXIF metadata following the formulation in [19]. (c) Real paired data without EXIF. For real pairs lacking EXIF, we obtain $S_{1}$ as in (b), and follow Eq. (2) to estimate the bokeh level $K$.
@@ 3.1 Stage 1: Defocus Deblurring
DeblurNet. We condition the deblurring network on a pair $(I_{\text{in}},I_{\text{pd}})$, where $I_{\text{pd}}$ is a conservative pre-deblur generated by an off-the-shelf defocus deblurring model. In our implementation, we use a non-generative defocus model  [56], whose outputs are typically content-faithful yet over-smoothed. Our model, guided by a learned diffusion prior, then injects high-fidelity details without sacrificing geometric correctness.
Artifact-robust dual conditioning.
To prevent artifacts in $I_{\text{pd}}$ from biasing the reconstruction and to ensure that fine details are recovered from $I_{\text{in}}$, we introduce two complementary mechanisms (see Fig. 2 (a)):
  • Positional disentanglement. We encode the two inputs with disjoint 2D positional grids: $I_{\mathrm{in}}$ is anchored at $(0,0)$, while $I_{\mathrm{pd}}$ is shifted to start at $(0,N)$, where $N$ denotes the side length of the token grid after downsampling $I_{\mathrm{in}}$. This shift lets the network distinguish the two sources, using $I_{\mathrm{pd}}$ as a low-frequency anchor while extracting high-frequency evidence from $I_{\mathrm{in}}$.
  • Pre-deblur dropout. During training, we stochastically replace $I_{\mathrm{pd}}$ with a zero tensor (at a fixed probability), forcing the network to remain effective when the auxiliary input is unreliable or absent. This regularizes the fusion, shields against method-specific artifacts, and enables the model to operate in a single-input deblurring mode at test time.
@@ 3.2 Stage 2: Shallow Depth-of-Field Synthesis
Synthetic-pair learning. BokehNet expects an AIF image, a target bokeh image, and the corresponding defocus map (see Fig. 2 (b)). In practice, perfectly paired data and accurate defocus maps are hard to acquire. We therefore pretrain with synthetic pairs: starting from real all-in-focus images and their estimated depth, we randomly sample a focus plane $S_{1}$ and a target bokeh level $K$, and use a simulator [48] to render bokeh consistent with the induced defocus map (see Fig. 3 (a)). This stage quickly trains the network to modulate the circle of confusion based on the defocus map, but it remains constrained by renderer bias and may introduce non-realistic artifacts.
Real unpaired bokeh learning. To capture real optics, we further leverage unpaired real bokeh images with EXIF metadata (see Fig. 3 (b)) and estimate an approximate bokeh level K\!K following [19]:
@@
$K\;\approx\;\frac{f^{2}\,S_{1}}{\,2\,F\,(S_{1}-f)\,}\times\textit{pixel\_ratio},$ (1)
where $f$ is the focal length and $F$ denotes the aperture $f$-number; and pixel_ratio compensates for sensor size and image resolution differences. The AIF input is provided by DeblurNet. Although $S_{1}$ can sometimes be read from metadata, both estimated and true depths may deviate in real scenes. To obtain more reliable supervision, we estimate in-focus masks using BiRefNet [88], and have human experts filter erroneous cases and complete partial masks. These masks refine the focus plane and yield a consistent defocus map, enabling BokehNet to learn from real bokeh images.
Real paired bokeh images without metadata.
@@
Previous bokeh-synthesis datasets [60] [30] provide real pairs but omit EXIF metadata or provide insufficient fields to estimate a reliable $K$ (see Fig. 3 (c)). We therefore adopt a simulator-in-the-loop calibration of the bokeh level: given an AIF image $I_{\mathrm{aif}}$, an estimated depth $D$, and a focus annotation obtained as in the unpaired setting (BiRefNet-based in-focus masks with expert filtering), we sweep $K$ and choose the value whose rendered result best matches the real bokeh target $I_{\mathrm{real}}$:
@@
$K^{\star}=\operatorname*{argmax}_{\begin{subarray}{c}K\in\\ (K_{\min},\,K_{\max})\end{subarray}}\operatorname{SSIM}\!\big(B(I_{\mathrm{aif}},D;\,S_{1},K),\,I_{\mathrm{real}}\big),$ (2)
where B denotes our physically guided renderer and $S_{1}$ is the focus-plane proxy. The selected $K^{\star}$ is then used as the pseudo bokeh level label for training . Focus-plane annotation follows the same protocol as in the unpaired case.
Discussion.
Although the resulting defocus maps and the AIF proxy are not perfectly accurate, this supervision regime mirrors the practical refocusing scenario. Aggregating numbers of real bokeh images substantially improves realism and robustness over synthetic-only training.

### v3

@@ Abstract
Depth-of-field control is essential in photography, but achieving perfect focus often requires multiple attempts or specialized equipment. Single-image refocusing is still difficult. It involves recovering sharp content and creating realistic bokeh. Current methods have significant drawbacks. They require all-in-focus inputs, rely on synthetic data from simulators, and have limited control over the aperture. We introduce Generative Refocusing, a two-step process that uses DeblurNet to recover all-in-focus images from diverse inputs and BokehNet to create controllable bokeh. This method combines synthetic and real bokeh images to achieve precise control while preserving authentic optical characteristics. Our experiments show we achieve top performance in defocus deblurring, bokeh synthesis, and refocusing benchmarks. Additionally, our Generative Refocusing allows custom aperture shapes. Project page: https://generative-refocusing.github.io/
Figure 1: Generative refocusing with controllable bokeh. Our model turns a single input image into a virtual camera that users can control. This enables a range of adjustments after image capture. (a) demonstrates aperture size control, allowing the user to change the depth of field from strong bokeh to an all-in-focus image. (b) demonstrates focus plane control by shifting the sharp region from the middle subject to the background. (c) highlights aperture-shape control, synthesizing a creative heart-shaped bokeh from point lights in the scene. (d) shows composite control, where both the focus plane and aperture size are adjusted simultaneously to reframe the subject.
@@ 1 Introduction
Controlling depth of field and focus is essential for artistic photography. It helps guide the viewer’s eye through selective focus and bokeh effects. Achieving perfect focus often requires multiple shots or specialized equipment, which can be challenging for everyday photographers. Single-image refocusing removes these obstacles. It allows for focus and bokeh adjustments after the photo is taken. The main challenge is recovering sharp details from blurred areas while creating natural-looking bokeh, while still giving users fine control.
@@
Existing research has addressed various aspects of this problem. Defocus deblurring methods [1, 8, 32, 53, 65, 71, 82] aim to recover sharp images. Diffusion approaches [36, 37] show promise for addressing spatially variant blur. Bokeh synthesis methods [9, 27, 43, 57] focus on rendering realistic depth-of-field. Recent efforts include physics-based [34, 33, 45], neural [43, 44], and diffusion-based [19, 80, 69] approaches. However, most methods handle either deblurring or bokeh synthesis individually. End-to-end refocusing often requires specialized capture setups [35, 41, 2].
@@
Despite these advances, three main limitations remain (Tab.˜1). First, most methods rely on all-in-focus inputs and accurate depth maps. This restricts their use to images that already have defocus blur. Second, while synthetic data [43, 87, 69] can be used for training, its realism is limited by the fidelity of the simulators. This approach often fails to capture detailed visual appearances. On the other hand, real datasets [57, 64] suffer from a critical limitation: they are restricted to isolated one-dimensional variations (i.e., solely aperture sweeps or focus plane sweeps). Consequently, they fail to provide the joint distribution of focus and aperture changes required to train a model for flexible, dual-parameter refocusing. Third, current methods typically support only aperture size, not its shape. This restriction limits the possibilities for creative bokeh effects.
@@
Table 1: Comparison across methods.  = supported;  = not supported; $\triangle$} = supported with caveat. Flexible input indicates compatibility with both all-in-focus (AIF) and defocused images. BokehMe and Bokehdiff require AIF inputs, while Learn2refocus strictly uses defocused inputs. For BokehMe, only the classical renderer variant supports aperture-shape control. Bokehlicious typically assumes an all-in-focus input and requires fine-tuning for different input types. Learn2Refocus generates focal stacks at discrete depth planes rather than supporting point-based focus selection.
@@
Method Flexible [1pt]input? Synthetic training data Real training data Supported user control types
@@
Focus- [1pt]sweep Aperture- [-1pt]sweep Single [-3pt]images Aperture [-3pt]size Aperture [-3pt]shape Focus [-1pt]point Defocus map
BokehMe [43] $\triangle$}
Bokehlicious [57]
BokehDiff [87]
DiffCamera [69]
Learn2Refocus [64] $\triangle$}
GenRefocus (Ours)
@@
To tackle these issues, we present Generative Refocusing (GenRefocus), a flexible single-image refocusing system built on a two-stage design. Our DeblurNet module produces a sharp, in-focus image from a variety of inputs. It uses a diffusion model guided by initial deblurring predictions. Our BokehNet generates fully customizable bokeh, accounting for user-defined focus planes, bokeh intensity, and aperture shapes (Fig.˜1). A significant advancement of our method is the proposed training scheme. While we utilize synthetic data to maintain geometric consistency, our key contribution lies in handling real-world data. By leveraging auxiliary information and dedicated processing, we complete the missing input or control signals (i.e., defocus maps) in existing real-world datasets. This strategy enables our model to effectively learn authentic lens characteristics that simulators fail to capture. Our approach achieves top performance in defocus deblurring, bokeh generation, and refocusing, while maintaining the natural consistency of scenes.
@@
  • A flexible refocusing pipeline accepts images in any focus state. It provides users with control over the focus plane, bokeh intensity, and aperture shape via two-stage decomposition.
  • A novel training scheme that overcomes the limitations of existing real-world datasets by completing absent input or control signals. This strategy successfully unites the geometric consistency of synthetic data with the authentic aberrations of real lenses.
  • The system performs well across defocus deblurring, bokeh synthesis, and refocusing. This has been validated on existing benchmarks (DPDD [1], RealDOF [32]) and our new light-field datasets (LF-Bokeh, LF-Refocus). It also has applications in creating aperture shapes.
@@ 2 Related Work
Figure 2: Pipeline Overview. Our method decomposes single-image refocusing into two stages: (a) Defocus Deblurring and (b) Shallow Depth-of-Field Synthesis. (a) The noisy latent $X_{t}$ and the encoded blurry input $I_{\text{in}}$ are concatenated into a unified token sequence $\mathbf{S}_{t}$. DeblurNet iteratively denoises this sequence to reconstruct a high-quality all-in-focus image $I_{\text{aif}}$. (b) Following the same unified-sequence design, BokehNet conditions on $I_{\text{aif}}$ and the defocus map $D_{\text{def}}$ to synthesize the refocused output $I_{\text{out}}$, with an optional aperture-shape condition AA. The VAE encoder $\mathcal{E}$ and decoder $\mathcal{D}$ map images to latent representations for the DiT backbone. The defocus map $D_{\text{def}}$ is computed from the estimated depth map $D$ [7] together with the user-specified focus plane $D_{\text{focus}}$ and bokeh level $K$.
2.0.1 Diffusion Models for Image Restoration.
Diffusion models [22, 61, 79, 66] advance image restoration via generative priors, extending to zero-shot restoration and artifact removal. Methods evolved from pixel-space [54] to efficient latent-space approaches [51, 74], enabling faster inference. ResShift [81] cuts steps from 1000 to 4–15 via residual shifting. Deblurring has progressed from merging diffusion priors with regression tasks, such as matting [72] or hierarchical restoration [12], to learning spatially varying kernels in latent space [30]. Cascaded pipelines [23] refine at multiple resolutions, while principled frameworks [40, 39] treat degradations as stochastic processes. Training-free methods [88, 28] enable plug-and-play restoration. Unlike these general methods, we target spatially-varying defocus blur with a two-stage pipeline explicitly separating deblurring from bokeh synthesis for controllable refocusing.
2.0.2 Defocus Deblurring.
@@
2.0.3 Bokeh Rendering.
Bokeh rendering progressed from physically based scattering [45, 34, 33] and differentiable rendering [60] to neural methods combining ray tracing with learned enhancements [43, 44]. Learning-based approaches evolved from fixed apertures [27] to variable f-stops [57, 9], extending to video [78] and 3D scenes [59, 70, 26]. Diffusion models enabled bokeh control [19, 80], mainly for text-to-image synthesis. These methods require all-in-focus inputs and accurate depth. We relax this constraint via flexible input handling, which accepts defocused or fully in-focus images.
2.0.4 Single-Image Refocusing.
Post-capture refocusing evolved from deconvolution [4, 84] and GANs [55] to recent diffusion models [46], yet significant limitations remain (see Tab.˜1). Constrained by conditioning on explicit focus coordinates, DiffCamera [69] requires fixed-resolution inputs ($512\times 512$) and suffers performance drops on arbitrary aspect ratios. Furthermore, it inherits simulator artifacts from its purely synthetic training. Learn2Refocus [64] trains on real data but lacks aperture-size control and generates a focal stack at predefined depth planes that may not align with the user’s target focus. Additionally, 3D representations [59, 70, 31, 38] offer controllable DoF but require multi-view capture. To address these gaps, we combine three capabilities: (1) Hybrid training scheme mixing synthetic and real data for authentic optics, (2) single-image flexibility accepting any input without preprocessing, and (3) comprehensive control over focus, intensity, and aperture shape.
2.0.5 Camera-Conditioned Diffusion.
Conditioning diffusion models on camera parameters progressed from extrinsic control for view synthesis [24, 13] and video generation [21, 73], to intrinsic control of focal length, aperture, and exposure [18, 11]. Direct conditioning often yields inconsistent content. Recent work uses temporal modeling [80], explicit calibration [6, 15], or transformer-based pose handling [3, 56]. These focus on generation; we operate in editing, manipulating images with scene consistency. By decoupling deblurring and bokeh stages, we achieve efficient single-image manipulation without multi-frame overhead.
@@ 3 Method
Single-image refocusing couples two operations: recovering sharp content in out-of-focus regions and applying controllable bokeh to originally sharp areas. Blur magnitude is dictated by a defocus map parameterized by a user-specified focus point and the scene depth. When the input is blurry, monocular depth estimation is brittle, undermining precise control over the defocus map and often yielding misfocus or artifacts. We propose Generative Refocusing (GenRefocus), which decomposes the task into defocus deblurring and bokeh synthesis, and comprises two models, DeblurNet and BokehNet, that jointly enable precise and controllable refocusing (see Fig. 2 for an overview of the pipeline); Such a flexible pipeline allows us to exert greater control over input conditions (see Tab. 1).
Figure 3: Training data generation. Each training sample consists of five components: (i) a bokeh image, (ii) an all-in-focus (AIF) image, (iii) a depth map $D$, (iv) a bokeh level $K$, and (v) a focus plane $D_{\text{focus}}$. We construct these samples via three routes: (a) Synthetic data. Given real AIF images and depth maps $D$, we compute a defocus map $D_{\text{def}}$ parameterized by randomly selected bokeh level $K$ and focus plane $D_{\text{focus}}$, and feed it into a bokeh renderer [43] to synthesize corresponding bokeh images. (b) ITW dataset [19]. Given real bokeh images, DeblurNet recovers an AIF image. We then estimate depth and extract a foreground mask [86] to define the estimated focus plane $D_{\text{focus}}$. The bokeh level $K$ is computed from the EXIF metadata and the estimated $D_{\text{focus}}$ following the formulation in [19]. (c) LFDOF [52] and RealBokeh. For real pairs, we obtain $D_{\text{focus}}$ as in (b), and follow Eq. (2) to estimate the bokeh level $K$.
@@ 3.1 Stage 1: Defocus Deblurring
Given a blurry input image $I_{\text{in}}$, we follow the conditioning design of [62, 63]. At each denoising step $t$, DeblurNet processes a single unified token sequence
$\small\mathbf{S}_{t}=\bigl[\,X_{t}\,;\,\mathcal{E}(I_{\text{in}})\,\bigr],$ (1)
where $X_{t}$ denotes the noisy image latent tokens at timestep $t$, $\mathcal{E}$ is a pretrained VAE encoder that maps the input image to latent condition tokens, and $\cdot\,$$\cdot\,$] denotes token concatenation. After iterative denoising, we decode the image latent part with the VAE decoder $\mathcal{D}$ to obtain the all-in-focus output $I_{\text{aif}}$ (Fig. 2 (a)). We train DeblurNet in a supervised manner using real paired data [1, 57].
@@ 3.2 Stage 2: Shallow Depth-of-Field Synthesis
Given an all-in-focus image $I_{\text{aif}}$ and a user-specified focus plane $D_{\text{focus}}$ with bokeh level $K$, we define the defocus map $D_{\text{def}}$ as
$\small D_{\text{def}}=K\cdot\left|D-D_{\text{focus}}\right|,$ (2)
where $D$ is the monocular depth map estimated from $I_{\text{aif}}$ using an off-the-shelf depth estimator [7]. Following the unified-token conditioning design in Stage 1, BokehNet concatenates the condition tokens of $I_{\text{aif}}$ and $D_{\text{def}}$ with the noisy latent tokens at each denoising step. BokehNet then iteratively denoises $X_{t}$ to synthesize the refocused output $I_{\text{out}}$, as illustrated in Fig. 2 (b).
3.2.1 Training data and supervision.
Training BokehNet requires $(I_{\text{aif}},I_{\text{out}},D_{\text{def}})$. Because accurately paired real supervision is scarce, we construct supervision via three complementary sources (Fig. 3): (a) synthetic data, (b) ITW dataset[19], and (c) LFDOF[52] and RealBokeh[57].
(a) Synthetic data.
@@
We pretrain with synthetic data: starting from real all-in-focus images and their estimated depth map $D$, we randomly sample a focus plane $D_{\text{focus}}$ and a target bokeh level $K$ (Fig. 3 (a)). We then construct $D_{\text{def}}$ using Eq. 2 and use a simulator [43] to render the corresponding target bokeh image consistent with $D_{\text{def}}$. This synthetic pretraining helps the network modulate the circle of confusion according to $D_{\text{def}}$; however, it is constrained by renderer bias and may introduce unrealistic artifacts.
(b) ITW dataset.
To capture real optics, we leverage ITW dataset[19] contains real bokeh images (see Fig. 3 (b)). For each real image, we use DeblurNet to produce the AIF input and compute an approximate bokeh level K\!K following [19]:
@@
$\small K\;\approx\;\frac{f^{2}\,D_{\text{focus}}}{\,2\,F\,(D_{\text{focus}}-f)\,}\times\text{pixel\_ratio},$ (3)
where $f$ is the focal length and $F$ is the aperture $f$-number, both directly obtained from EXIF metadata. The pixel_ratio term accounts for differences in camera sensor size and image resolution. Although some devices may provide a focus-distance field in EXIF, it is frequently missing or noisy; therefore, we do not rely on EXIF for $D_{\text{focus}}$. Instead, we compute the focus plane based on an in-focus mask and a monocular depth estimate, following the approach of  [19]. Specifically, we estimate an in-focus mask $M$ using BiRefNet [86] and produce a depth map $D$ using a monocular depth estimator. The focus plane is then determined as the median depth within the masked in-focus area:
$\small D_{\text{focus}}=\mathrm{median}\bigl(D[M]\bigr).$ (4)
We then generate $D_{\text{def}}$ using Eq. 2, enabling BokehNet to effectively train on real single images. While these estimated AIF images and defocus maps contain inevitable noise, training on them inherently aligns with our single-image refocusing pipeline, making the model more robust to depth and deblurring artifacts.
(c) LFDOF and RealBokeh.
These datasets provide pairs but omit EXIF metadata or provide insufficient fields to estimate $K$ (see Fig. 3 (c)). To address this, we first compute the focus-plane proxy $D_{\text{focus}}$. Similar to (b), we employ BiRefNet [86] to obtain an initial in-focus mask $M$. However, due to the increased diversity and complexity of the scenes in these datasets, the initial estimate of $M$ is sometimes unreliable. Rather than simply verifying and discarding unreliable cases, we introduce a manual refinement step. Specifically, we re-select a small yet reliable in-focus region to correct $M$, thereby accurately extracting $D_{\text{focus}}$. This strategy preserves challenging samples rather than excluding them, thereby enabling the model to learn from a wider range of real-world scenarios.
@@
With $D_{\text{focus}}$ established, we adopt a simulator-in-the-loop calibration of the bokeh level. Given an AIF image $I_{\mathrm{aif}}$ and estimated depth $D$, we sweep $K$ and choose the value whose rendered result best matches the real bokeh target $I_{\mathrm{real}}$:
@@
$\small K^{\star}=\operatorname*{argmax}_{\begin{subarray}{c}K\in\\ (K_{\min},\,K_{\max})\end{subarray}}\operatorname{SSIM}\!\big(R(I_{\mathrm{aif}},D;\,D_{\text{focus}},K),\,I_{\mathrm{real}}\big),$ (5)
where $R$ denotes our physically guided renderer. The selected $K^{\star}$ is then used as the pseudo-bokeh-level label for training, provided that its corresponding SSIM exceeds a predefined threshold to ensure reliable supervision. Finally, the defocus map $D_{\text{def}}$ is constructed accordingly following Eq. 2.